3x^2-2y^2+(1-4xy)y'=0

1
Expert's answer
2021-11-24T18:22:13-0500

Let us solve the differential equation "3x^2-2y^2+(1-4xy)y'=0" which is equivalent to the equation "(3x^2-2y^2)dx+(1-4xy)dy=0." Since "\frac{\partial(3x^2-2y^2)}{\partial y}=-4y=\frac{\partial(1-4xy)}{\partial x}," we conclude that this equation is exact, and hence "U_x=3x^2-2y^2, U_y=1-4xy" for some function "U=U(x,y)." It follows that "U=x^3-2y^2x+C(y)," and hence "\frac{\partial U}{\partial y}=-4yx+C'(y)." Therefore, "-4yx+C'(y)=1-4xy," and hence "C(y)=y+C."

We conclude that the general solution of the differential equation is

"x^3-2y^2x+y=C."